OpenStudy (anonymous):

what's -4(4+7y)+5+3y

OpenStudy (anonymous):

Try using P.E.M.D.A.S.

OpenStudy (hockeychick23):

Use Order of operations or PEMDAS Once you use the distributive property you will get this -16-28y+5+3y Now you need to simplify -28y+3y= -25y 5-16= -11 -25y-11